Find the equation of the straight line passing through the point( 1, 2, 3 )to intersect the straight line X + 1 = 2(y-2)=z+4 and parallel to the plane X + 5 y + 4 z = 0

Grade:

1 Answers

Vikas TU
5 years ago
Equation of the straight line passing through the point ( 1, 2, 3 ):
(x-1)/1 = (y-2)/5 = (z-3)/4
Let 
(x-1)/1 = (y-2)/5 = (z-3)/4 = k
then, x = k+1
y = 5k + 2
z =4k + 3
putting all the 3 points in the straight line equation we get,
k+1 = 2(5k + 2 – 2)
10k = k + 1
9k = 1
k = 1/9
Hnece the intersecting points are:
x = 11/9
y = 23/9
z = 31/9.